OpenStudy (anonymous):

Write this equation in the standard form of a hyperbola and identify the vertices. x^2 − 4y^2 = 100

OpenStudy (anonymous):

divide by 100 for standard form\[\frac{x^2}{100} − \frac{4y^2}{100} = \frac{100}{100}\]simplify\[\frac{x^2}{100}-\frac{y^2}{25}=1\]Center (0,0) Transverse axis y=0 a=10 => Vertices @ (-10,0) and (10,0)